PROBLEM 2-14/Classification of Labour Costs ILO1] Big Sky Equipment manufact ures and sells camping equipment. Employees in the production department are paid $30 per hour but receive $45 per hour for each hour worked in excess of 40 hours per week. Required: 1. Calculate the total wages for the week if an employee works 50 hours. Of the total, how much would the company allocate to direct labour cost? To manufacturing overhead cost? In a different week, suppose an employee works 45 hours but is idle for 3 hours during the week due to a shortage of direct materials used in production. Calculate the employ- ees total wages for the week. How much of this amount would be allocated to direct la- bour cost? To manufacturing overhead cost? Big Sky provides employee benefits that cost $9 for each hour the employee works (either regular time or overtime). During a particular week, an employee works 52 hours but is idle for 6 hours due to an equipment breakdown. Calculate the employees total wages and employee benefits for the week. If employee benefits are considered by the company to be part of manufacturing overhead cost, how much of the total wages and employee benefits for the week would be allocated to direct labour cost? To manufacturing overhead cost? Refer to the data in (3) above. If the company treats that part of employee benefits relating to direct labour as added direct labour cost, how much of the wages and employee bene- fits for the week will be allocated to direct labour cost? To manufacturing overhead cost? 2. 3. 4.
Answer #1
1 Total hours worked 50 hours
Wages rate upto 40 hrs $30
Wages rate after 40 hrs $45
Total wages=
for 40 hours 1200
after 40 hours i.e. 50-40=10 hours= 450
Total wages= 1650
Direct labour cost=(40*30) 1200
Manufacturing overhead cost=(10*45) 450
2 Total hours worked 45 hours
Wages rate upto 40 hrs $30
Wages rate after 40 hrs $45
Total wages=
for 40 hours 1200
after 40 hours i.e. 45-40=5 hours= 225
Total wages= 1425
Direct labour cost=(40*30+3*45) 1335
Manufacturing overhead cost=(2*45) 90
3 Total hours worked 52 hours
Wages rate upto 40 hrs $30
Wages rate after 40 hrs $45
Employee benefits rate $9
Total wages=
for 40 hours 1200
after 40 hours i.e. 52-40=12 hours= 540
Total wages= 1740
Total employee benefits 468
Direct labour cost=(40*30+9*45) 1605
Manufacturing overhead cost=(3*45)+52*9 603
4 Total hours worked 52 hours
Wages rate upto 40 hrs $30
Wages rate after 40 hrs $45
Employee benefits rate $9
Total wages=
for 40 hours 1200
after 40 hours i.e. 52-40=12 hours= 540
Total wages= 1740
Total employee benefits 468
Direct labour cost=(40*30+9*45)+(49*9) 2046
Manufacturing overhead cost=(3*45)+3*9 162